# SMC++: Masked Learning of Unsupervised Video Semantic Compression

### v1

@@ 3.4 Blueprint-Guided Compression Transformer
In this section, we introduce another new component of the SMC++ model, i.e., Blueprint-guided compression Transformer (Blue-Tr), which effectively captures redundancy among heterogeneous features across different frames and coding layers. As illustrated in Figure 3, Blue-Tr operates under an ”align-then-compress” paradigm, guided by a novel blueprint semantic representation. The compression procedure consists of the following three steps:
@@ 4.1 Evaluation Datasets
For action recognition task, we evaluate on four large-scale video datasets, UCF101 [119], HMDB51 [120], Kinetics [83], and Diving48 [121]. For multiple object tracking (MOT) task, we evaluate on MOT17 [122]. For video object segmentation (VOS) task, we evaluate on DAVIS2017 [123]. We also compare the compression performance on Standard HEVC Test Sequences [1], including HEVC Class B/C/D/E datasets.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Semantic coding performance on Action Recognition (1st and 2nd rows), VOS (3rd row), and MOT (4th row) tasks. The plot titles are in {Dataset}-{Task Model} format. The codec setting is LDP mode with GOP size 10.
Dataset Processing. During the training procedure, we randomly select 60K videos of resolution larger than 1280$\times$720 from the training set of Kinetics400, following the previous approach [37]. The quantization parameter (QP) value of VVC is randomly sampled, so that the framework is learned to adapt to various QPs with a single model. To reduce the training time cost, we pre-compress the training videos with various QP and buffer these compressed videos in an offline manner, so that the time-consuming video compression procedure can be eliminated.
For the evaluation of action recognition datasets, we first downsample the shortest side of the video to 256 pixels and then crop it to the size 224$\times$224 before the coding procedure. For the evaluation of MOT, we adopt the original MOT17 dataset without downsampling, since many tracking methods require high-resolution inputs. For the evaluation of VOS, we download the high-resolution version of DAVIS2017, which contains videos from 720p to 1080p, and then downsample them to 480p (854$\times$480), which is the input resolution of most VOS methods.
4.2 Experimental Setting
Downstream Task Models. For the action recognition task, we adopt the following popular models, i.e., TSM [124], SlowFast [125], and TimeSformer [126], including 2D CNN, 3D CNN and the recent Transformer architectures. The model weights are provided by the MMAction2 framework [127]. For the MOT task, we adopt ByteTrack [128], of which the model weights are provided by MMTracking [129]. For the VOS task, we adopt XMem [130], of which the model weights are officially released by the authors. We emphasize that we directly feed the decoded videos by our framework into these official models, without fine-tuning them, since fine-tuning the numerous downstream video models for the codec is not practical in real-world applications. The adopted models are right the same as VCS [37], as we prefer a completely fair comparison.
Baseline Codecs. HEVC is evaluated with FFmpeg software [131] and x265 codec. VVC is evaluated with VVenC1.5.0 software [132], which consumes much less encoding time than the VVC reference software VTM, so that supporting the evaluation of large-scale video datasets containing thousands of videos. FVC [3] is a representative work on learnable codec. PLVC [53] is a recent perceptual quality-oriented codec. DCVC-FM [38] is the state-of-the-art neural video codec. VCS [37] is the state-of-the-art semantic video compression method. For traditional codecs, the QP value is selected from {51,47,43,39,35}, as setting QP to larger values does not degrade the downstream task obviously. All codecs and our framework are evaluated on Low-Delay with P frame (LDP) coding mode. Both VCS and our models SMC and SMC++ are built upon the VVenC1.5.0 software, for a fair comparison.
Evaluation Metrics. We use bpp (bit per pixel) to measure the average number of bits used for one pixel in each frame. For the action recognition task, we adopt the Top1 accuracy as the performance indicator. For the MOT task, we adopt MOTA (multiple object tracking accuracy) [133], MOTP (multiple object tracking precision), FN (false negative detection number), and IDF1, which is the ratio of correctly identified detections over the average number of ground truth and computed detections. For the VOS task, the standard metrics Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$ and the average of $\mathcal{J}$ and $\mathcal{F}$ ($\mathcal{J\&F}$) are adopted. We also report the contour recall $\mathcal{F}$-$Recal$.
Implementation Details. Following VideoMAE [32], the masking ratio for the MAE task is set to 90%, where the encoder and decoder networks consist of six and two ViT blocks [134] with divided space-Time attention [126], respectively. Following previous works [135][84] on video temporal modeling, we also insert temporal convolutions with kernel size 3 before each 3$\times$3 convolutions of bottleneck layers of the original ResNet18 encoder network, aiming to enhance its basic temporal modeling ability. The temporal convolutions are implemented using causal convolutions [136]. In this design, only the historical information is accessible, with the intention of preserving the P-frame coding mode setting. The weights of all networks within our framework are randomly initialized.
$K$ in Equation 3 is empirically set as 5. $\alpha$ and $\beta$ are set to 1 and 0.1, respectively. The resolution of training videos is 256 $\times$ 256 and the clip length is 8. We use the Adam optimizer [137] by setting the learning rate as $1\times 10^{-4}$, $\beta_{1}$ as 0.9 and $\beta_{2}$ as 0.999, respectively. The learning rate is decayed to $1\times 10^{-5}$ in the last 100k iterations. The training iteration number of SMC is 1000k. The mini-batch size is 24. To ensure that the encoder begins with a well-initialized state, we omit the quantization step for the initial 100k iterations. This allows precise gradient values to be propagated back to the encoder side. The system is implemented with Pytorch [138] and it takes about six days to train the SMC model using eight Nvidia 3090Ti GPUs. The SMC++ model is initialized from the SMC model and trained by another 500,000 iterations. The learning rate is set to $1\times 10^{-4}$ in the first 400,000 iterations and decayed to $1\times 10^{-5}$ in the last 100,000 iterations. The additional training procedure of SMC++ takes about six days with eight Nvidia 3090Ti GPUs.
4.3 Experimental Results
@@
Metric MOTA(%) IDF1(%) MOTA(%) IDF1(%) MOTA(%) IDF1(%)
@@
Metric $\mathcal{J}$$&\mathcal{F}$ (%) $\mathcal{F}$ (%) $\mathcal{J}$$&\mathcal{F}$ (%) $\mathcal{F}$ (%) $\mathcal{J}$$&\mathcal{F}$ (%) $\mathcal{F}$ (%)
@@
[…]
@@
Action Recognition. In Table I, we evaluate various video compression methods on the action recognition task. Our basic SMC model remarkably outperforms traditional codecs HEVC and VVC, the learnable codecs FVC and DCVC-FM, and the perceptual codec PLVC. For example, on the UCF101-TSM setting, our SMC model shows 22% and 10% Top1 accuracy improvement over FVC and VVC, respectively, at 0.04bpp. On the HMDB51-TSM setting, SMC outperforms the DCVC-FM by 18% at 0.03bpp. Although the most recent semantic coding approach VCS surpasses our basic SMC model, our enhanced SMC++ model outperforms it again, notably with over 5% Top1 accuracy gain on the Diving48-TSM setting at 0.03bpp. Moreover, on the large-scale Kinetics400 dataset, our SMC++ model surpasses VCS and DCVC-FM by 1.86% and 13.42% Top1 accuracy, respectively, at 0.06bpp. When being evaluated with the Slowfast and TimeSformer action recognition networks, our SMC++ model still outperforms all other approaches by a large margin, i.e., about 4% Top1 gain over VCS at 0.02bpp.
We present the rate-performance (RP) curves of recent coding methods in the 1st and 2nd rows of Figure 4. Our SMC++ model leads all other methods across all datasets and bitrate levels. Our basic SMC model also excels, outperforming all except VCS. Interestingly, we observe that the state-of-the-art neural codec DCVC-FM shows unsatisfactory performances, e.g., it even underperforms VVC on the HMDB51-TSM and Kinetics400-TSM settings, despite having superior PSNR performance [38] compared to the successor of VVC, i.e., the ECM software [139]. This suggests that the recent neural codecs may be overfitting to the standard HEVC test sequences, as its hyperparameters are manually tuned for these sequences.
@@
Multiple Object Tracking (MOT). Beyond action recognition, we evaluated our coding approach on the more challenging Multiple Object Tracking (MOT) task, demanding precise localization and robust feature extraction under occlusions. Results are shown in Table II. At low bitrates like 0.01bpp, our SMC model outperforms FVC, PLVC, and VVC by 26.50%, 2.97%, and 5.98% in MOTA, respectively, matching DCVC-FM. Our SMC++ surpasses DCVC-FM by a large margin of 1.66%. Compared to the semantic coding approach VCS, SMC and SMC++ show a 3.05% and 4.72% MOTA gain, respectively.
At higher bitrates like 0.02bpp, DCVC-FM demonstrates strong performances, probably due to its internal high-resolution feature designs. Despite this, SMC++ closely matches DCVC-FM in MOTA (75.33% vs. 75.58%) and exceeds it in IDF1 (76.90% vs. 75.56%).
We also notice that, although FVC performs much better than the legacy HEVC codec on the action recognition task, it performs surprisingly poor on the MOT task, with a significantly lower MOTA (44.24% v.s. 61.30% of HEVC) at the 0.01bpp. This is due to FVC’s reliance on the hand-crafted BPG codec for I frames and a neural codec for P frames, causing appearance inconsistency and hindering tracking consistency. DCVC-FM mitigates this through end-to-end optimization of I- and P-frame compression networks.
@@
Finally, we provide the RP curves. As shown in the third row of Figure 4, SMC outperforms all previous methods except DCVC-FM, while SMC++ leads across most metrics. At higher bitrates, SMC++’s MOTP performance is unsatisfactory, which denotes the bounding-box accuracy of the detected objects. We conjecture the reasons is that the Transformer-based Blue-Tr module within SMC++ emphasizes high-level features and compromises local positioning accuracy. We plan to fix this by adaptively choosing between CNN- and Transformer-based compression modules.
@@
Video Object Segmentation (VOS). We further evaluate different coding methods on the VOS task. As shown in Table III, our basic SMC model surpasses all previous approaches, except VCS and DCVC-FM. For example, SMC exceeds VVC and PLVC by 6.73% and 12.75% in terms of $\mathcal{J}$$&\mathcal{F}$ at the 0.01bpp level. After introducing the masked motion prediction and Blue-Tr module, our resultant SMC++ model outperforms VCS by 3.04% and 3.68% in terms of $\mathcal{J}$$&\mathcal{F}$ and $\mathcal{F}$ respectively, and DCVC-FM by 1.30% and 3.21% in the same metrics, at 0.01bpp. The RP curves of recent approaches are provided in the fourth row of Figure 4.
@@
More Advanced Test Conditions. Recent learnable codecs [4][140][48] have adopted a more challenging test condition, i.e., group-of-picture (GOP) size of 32. This larger GOP size poses a greater challenge than small GOP sizes, because of the error propagation issue. As shown in Figure 5, our advanced SMC++ model maintains its effectiveness under this more advanced test condition.
In the context of action recognition, SMC++ achieves approximately 2%, 5%, and 2% higher Top-1 accuracy compared to VVC, DCVC-FM, and VCS, respectively, on the HDMB51-SlowFast setting at 02bpp. Likewise, in the VOS task, SMC++ outperforms VVC, DCVC-FM, and VCS by about 5%, 2% and 2.2% in terms of the $\mathcal{J\&F}$ metric, respectively, on the DAVIS2017-XMem setting at 0.01bpp.
Furthermore, it’s worth noting that the basic SMC model performs significantly worse than the VVC codec on the VOS task, despite being built upon and intended to enhance the VVC codec. For instance, at 0.01bpp, the VVC codec attains a $\mathcal{J\&F}$ score of approximately 75%, whereas SMC manages only 70%. This discrepancy arises from SMC’s simple abstraction-based compression module, which struggles to handle the feature domain shift encountered in large GOP scenarios. In contrast, the SMC++ model boosts VVC’s performance by 5% $\mathcal{J\&F}$ at 0.01bpp. This proves the Blueprint-based feature alignment operations in SMC++ effectively mitigates the error propagation problem.
@@
Visual Quality Assessment. As shown in Table IV, our models SMC and SMC++ substantially outperform the previous traditional VVC, neural codec DCVC-FM, as well as semantic coding approach VCS in terms of perceptual visual quality, as evidenced by the much lower LPIPS values. We further demonstrate the frames compressed by different approaches. As shown in Figure 6, our approach shows sharper edges and better object structures than VVC and DCVC-FM, while showing fewer artifacts than VCS.
TABLE IV: Comparison of different methods on HEVC Class C. The results are reports in PSNR/LPIPS format. A lower LPIPS value indicates better perceptual quality.
0.01bpp 0.02bpp 0.03bpp
VVC 24.32/0.4824 26.21/0.4104 27.46/0.3690
DCVC-FM 25.02/0.4927 27.11/0.4170 28.26/0.3778
VCS 20.86/0.5365 23.33/0.4062 23.81/0.3683
SMC 20.96/0.4083 23.29/0.3307 23.98/0.3058
SMC++ 20.92/0.3936 23.92/0.3129 24.75/0.2890
Figure 6: Qualitative comparison of different approaches.
Towards High-fidelity Compression. Despite offering superior visual pleasure in perceptual quality, our approach yields lower PSNR results compared to VVC and DCVC-FM. For instance, as shown in Table IV, our SMC++ model lags behind DCVC-FM by roughly 3.5dB at the 0.03bpp level. Given the importance of high-fidelity decoding for the secondary manual review of AI analysis results, we’ve enhanced our SMC++ model by appending a DR-Net, creating the SMC++* model. This aims to merge our decoded video with the lossy VVC-based video, enriching it with details. To match the performance of the state-of-the-art neural codec DCVC-FM [38], we utilize the advanced VTM20.0 software [39] as the base VVC coding layer. Since high-fidelity decoding usually requires a higher bitrate than semantic coding applications, we set the QP of the VTM20.0 software to larger values, i.e., 23, 26, 29, and 32.
@@
As shown in the green curves of Figure 7, when setting VTM20.0 to LDP mode, SMC++ notably improves performance on various datasets. For instance, on the HEVC class D dataset, we observe about 0.4dB PSNR gain at the 0.2bpp bitrate level. We also appended DR-Net to VCS and SMC models, training it for both, resulting in VCS* and SMC* models. On the HEVC class D dataset at the 0.2bpp bitrate level, VCS*, SMC*, and SMC++* achieved performances of 33.90dB, 33.82dB, and 34.01dB, respectively. This trend aligns with their semantic modeling capability, since good semantics is beneficial to video detail reconstruction [141].
Furthermore, when applied to the more advanced random access (RA) mode of VTM20.0, our SMC++ still demonstrates substantial improvements. As indicated by the red curves of Figure 7, we observe about 0.3dB quality enhancement at the 0.25bpp bitrate, on the HEVC class C dataset. We have observed that our method doesn’t enhance coding performance on the HEVC Class E dataset at low bitrates. This is likely because this dataset consists of conference videos with simplistic scenes and objects (humans and desks), lacking complex semantic elements like intricate object arrangements and interactions. As a result, the rich semantics learned by our approach aren’t particularly helpful in reconstructing these videos.
It seems like our non-semantic suppressed (NSS) learning strategy, excluding the non-semantic detail information from the coding procedure, is contradictory to the detail decoding task. However, our framework indeed improves the base coding layer VTM, since this prompts a “divide-and-conquer” paradigm, where the VTM layer and our semantic layer focus on coding textual information and semantic information, respectively. The semantic information such as scene structure and object shape effectively corrects the distorted textures of VTM.

### v2

@@ 3.5 Blueprint-Guided Compression Transformer
In this section, we introduce another new component of the SMC++ model, i.e., Blueprint-guided compression Transformer (Blue-Tr), which effectively captures redundancy among heterogeneous features across different frames and coding layers. As illustrated in Figure 2, Blue-Tr operates under an “align-then-compress” paradigm, guided by a novel blueprint semantic representation. The compression procedure consists of the following three steps:
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Semantic coding performance on Action Recognition (1st and 2nd rows), VOS (3rd row), and MOT (4th row) tasks. The plot titles are in {Dataset}-{Task Model} format. The codec setting is LDP mode with GOP size 10. The results on more advanced codec settings are provided in the supplementary material.
@@ 4.1 Experimental Details
Evaluation Datasets. For action recognition task, we evaluate on four large-scale video datasets, UCF101 [115], HMDB51 [116], Kinetics [84], and Diving48 [117]. For multiple object tracking (MOT) task, we evaluate on MOT17 [118]. For video object segmentation (VOS) task, we evaluate on DAVIS2017 [119]. We also compare the compression performance on Standard HEVC Test Sequences [1], including HEVC Class B/C/D/E datasets. Dataset processing pipeline is provided in the supplementary material.
Experimental Setting. For the downstream tasks, we use TSM [120], SlowFast [121], and TimeSformer [122] for action recognition, ByteTrack [123] for MOT, and XMem [124] for VOS; for baseline codecs, HEVC, VVC [125], FVC [3], PLVC [45], DCVC-FM [40], and VCS [72] are evaluated. All codecs and our framework are evaluated on Low-Delay with P frame (LDP) coding mode. The evaluation metrics include bpp, Top1 accuracy (action recognition), MOTA [126], MOTP, FN, and IDF1 (MOT), as well as the Jaccard index ($\mathcal{J}$), contour accuracy ($\mathcal{F}$), $\mathcal{J\&F}$, and $\mathcal{F}$-Recall (VOS). For more details, please refer to the supplementary material.
@@
Metric
MOTA(%)
IDF1(%)
MOTA(%)
IDF1(%)
MOTA(%)
IDF1(%)
@@
Metric $\mathcal{J}$$\mathcal{F}$ (%) $\mathcal{F}$ (%) $\mathcal{J}$$\mathcal{F}$ (%) $\mathcal{F}$ (%) $\mathcal{J}$$\mathcal{F}$ (%) $\mathcal{F}$ (%)
@@
gray[5pt/3pt] Original 87.70 91.33 87.70 91.33 87.70 91.33
Implementation Details. We set the masking ratio for the MAE task to 90%, and employ six encoder and two decoder ViT blocks with divided space-time attention [127, 122]. All network weights are randomly initialized. $K$, $\alpha$, and $\beta$ are set to 5, 1 and, 0.1, respectively. Training is performed on $\times256$ video clips (length 8) using the Adam optimizer [128] with an initial learning rate of $1\times 10^{-4}$, decayed to $1\times 10^{-5}$ during the final 100k iterations. The total iteration number is 1M. The mini-batch size is 24. The quantization step is omitted for the first 100k iterations to allow precise gradient propagation. The SMC model is trained in approximately six days, on eight Nvidia 3090Ti GPUs. SMC++ is subsequently initialized from SMC and further trained for 500k iterations, with the learning rate maintained at $1\times 10^{-4}$ for the first 400k iterations and decayed to $1\times 10^{-5}$ during the final 100k iterations, also taking about six days on eight Nvidia 3090Ti GPUs.
4.2 Experimental Results
@@
Action Recognition. In Table 4.1, we evaluate various video compression methods on the action recognition task. Our basic SMC model remarkably outperforms traditional codecs HEVC and VVC, the learnable codecs FVC and DCVC-FM, and the perceptual codec PLVC. For example, on the UCF101-TSM setting, our SMC model shows 22% and 10% Top1 accuracy improvement over FVC and VVC, respectively, at 0.04bpp. On the HMDB51-TSM setting, SMC outperforms the DCVC-FM by 18% at 0.03bpp. Although the most recent semantic coding approach VCS surpasses our basic SMC model, our enhanced SMC++ model outperforms it again, notably with over 5% Top1 accuracy gain on the Diving48-TSM setting at 0.03bpp. Moreover, on the large-scale Kinetics400 dataset, our SMC++ model surpasses VCS and DCVC-FM by 1.86% and 13.42% Top1 accuracy, respectively, at 0.06bpp. When being evaluated with the Slowfast and TimeSformer action recognition networks, our SMC++ model still outperforms all other approaches by a large margin, i.e., about 4% Top1 gain over VCS at 0.02bpp. The rate-performance (RP) curves of different methods are provided in the 1st and 2nd rows of Figure 4
TABLE IV: Comparison of different coding methods on FineGym99 [129], using TSM action model.
TSM Top1 (%)
Bpp @0.02bpp @0.04bpp @0.06bpp @0.08bpp
VVC 42.84 71.17 80.91 85.53
DCVC-FM 54.67 77.34 83.46 85.85
VCS 61.16 77.71 83.02 85.21
SMC 62.43 80.00 85.17 87.30
SMC++ 70.67 82.13 86.21 87.70
Finally, we evaluate the performance of our framework on the fine-grained action recognition dataset FineGym99 [129], which requires distinguishing subtle action differences in various gymnasium videos. As shown in Table IV, our SMC method already outperforms previous approaches by a large margin. For instance, at 0.02bpp, SMC achieves a Top-1 accuracy of 62.43%, whereas state-of-the-art learnable video codec DCVC-FM and semantic video codec VCS achieve only 54.67% and 61.16%, respectively. This demonstrates the strong fine-grained action discrimination capability of our approach. Our enhanced SMC++ further establishes a new state-of-the-art by building upon and improving SMC’s performance.
@@
Multiple Object Tracking (MOT). Beyond action recognition, we evaluated our coding approach on the more challenging Multiple Object Tracking (MOT) task, demanding precise localization and robust feature extraction under occlusions. Results are shown in Table 4.1. At low bitrates like 0.01bpp, our SMC model outperforms FVC, PLVC, and VVC by 26.50%, 2.97%, and 5.98% in MOTA, respectively, matching DCVC-FM. Our SMC++ surpasses DCVC-FM by a large margin of 1.66%. Compared to the semantic coding approach VCS, SMC and SMC++ show a 3.05% and 4.72% MOTA gain, respectively.
@@
We also provide the RP curves in Figure 4. SMC outperforms all previous methods except DCVC-FM, while SMC++ leads across most metrics. At higher bitrates, SMC++’s MOTP performance is unsatisfactory, which denotes the bounding-box accuracy of the detected objects. We conjecture the reasons is that the Transformer-based Blue-Tr module within SMC++ emphasizes high-level features and compromises local positioning accuracy. We plan to fix this by adaptively choosing between CNN- and Transformer-based compression modules.
@@
Video Object Segmentation (VOS). We further evaluate different coding methods on the VOS task. As shown in Table 4.1, our basic SMC model surpasses all previous approaches, except VCS and DCVC-FM. For example, SMC exceeds VVC and PLVC by 6.73% and 12.75% in terms of $\mathcal{J}$$\mathcal{F}$ at the 0.01bpp level. After introducing the masked motion prediction and Blue-Tr module, our resultant SMC++ model outperforms VCS by 3.04% and 3.68% in terms of $\mathcal{J}$$\mathcal{F}$ and $\mathcal{F}$ respectively, and DCVC-FM by 1.30% and 3.21% in the same metrics, at 0.01bpp. The RP curves of recent approaches are provided in the fourth row of Figure 4.
We further evaluate E-Nerv [130] on the VOS task, as the NerV [131] series has recently shown promising performance in video compression but has yet to be assessed on video understanding tasks. Due to the high computational cost of the current NerV [131] approaches, we tested only one sequence “bike-blackswan_1”, from the DAVIS2017 dataset. E-Nerv [130] achieves a $\mathcal{J}$ value of 82.28% @0.079 bpp, which is much inferior to our approach, i.e., $\mathcal{J}$ value of 90.28% at a even lower bitrate 0.048 bpp. This highlights the limitations of NerV approaches, as their overfitting nature restricts them to compressing low-level details. In contrast, our method, which leverages semantics learned from large-scale videos, achieves superior results.
Comparison to Other Recent Approaches. We compare our approach with two more recent methods: (1) DeepSVC [27], a scalable video semantic compression approach, and (2) HVFVC [132], a recent perceptual video compression method. As shown in Figure 5, HVFVC is much inferior to both SMC and SMC++ models, in both recognition and segmentation tasks. This is because HVFVC prioritizes perceptual quality but fails to effectively preserve semantic information critical for downstream tasks. At 0.025 bpp, HVFVC achieves only 77% accuracy, while SMC++ reaches 85% on UCF101-TSM, representing an 8% performance gap. DeepSVC is comparable to the baseline SMC model but falls notably short of SMC++. For instance, on UCF101-TSM at 0.075 bpp, DeepSVC, SMC, and SMC++ achieve 90.01%, 89.82%, and 91.41%, respectively. Similarly, on the VOS task, DeepSVC achieves about 81% $\mathcal{J}\&\mathcal{F}$ score, significantly lower than the 82.3% achieved by SMC++ on the DAVIS2017-XMEM setting at 0.03 bpp.
Refer to caption Refer to caption
Figure 5: Comparison of our approach with two recent approaches, i.e., DeepSVC [27] and HVFVC [132], on action recognition and video object segmentation (VOS) tasks.
Compatibility to Other Codecs as the Base Layer. Beyond using VVC as the base coding layer, we also evaluate our framework with the fast H.264 and H.265 codecs to assess improvements in coding efficiency, as well as with the state-of-the-art DCVC-FM to verify whether our proposed framework remains effective when applied to an already-strong codec. As shown in Figure 6, our approach consistently enhances performance across these various codecs on semantic tasks, action recognition, and VOS tasks.
Refer to caption Refer to caption
Figure 6: Performance enhancement of SMC++, when using H.264, H.265 and DCVC-FM as the base coding layer.
Refer to caption Refer to caption
Figure 7: Comparison of different approaches on VQA and VOS tasks, with LLM-based downstream models.
Effectiveness on LLM-based downstream models. Large multi-modal models are becoming increasingly popular for video analysis tasks. To evaluate the compatibility of our method with such models, we conducted additional experiments on the video question-answering (VQA) task using the MSVD-QA dataset [133] with MA-LMM [134] model, and on the video object segmentation (VOS) task using the DAVIS2017 dataset with PSALM [135] model.
As shown in Figure 7, for the VQA task, our approach demonstrates consistently strong performance. For example, at 0.02 bpp, our method outperforms VVC and DCVC-FM by 5.85% and 3.12%, respectively. This proves that our method effectively preserves semantic information crucial for cross-modality reasoning tasks, even at very low bitrates. For the VOS task, we evaluate with the LLM-based approach PSALM [135]. Our approach remarkably surpasses previous coding methods. For instance, at 0.01 bpp, our method exceeds VVC and DCVC-FM by 11.23% and 7.15%, respectively, in terms of the $\mathcal{J\&F}$ metric. This demonstrates that, even in the era of LLM, our video semantic compression strategy remains highly effective, supporting LLM-based video analysis models at low bitrate levels.
@@
More Advanced Test Conditions. Recent learnable codecs [4][39] have adopted a more challenging test condition, i.e., group-of-picture (GOP) size of 32. This larger GOP size poses a greater challenge than small GOP sizes, because of the error propagation issue. As shown in Figure 8, our advanced SMC++ model maintains its effectiveness under this more advanced test condition.
@@
Towards High-fidelity Compression. To enable high-fidelity decoding, we enhance SMC++ with DR-Net to form SMC++*, employing VTM20.0 as the base VVC layer. As shown in Figure 9, SMC++* achieves an approximately 0.3dB gain on the HEVC Class C dataset, compared to the advanced RA mode of VTM at 0.25bpp.